Example 4 - One card is drawn from a well-shuffled deck - Examples - Examples

part 2 - Example 4 - Examples - Serial order wise - Chapter 14 Class 10 Probability

Remove Ads Take short quiz All Quiz and Worksheets
Teachoo · Class 10 Explore Class 10

Transcript

Example 4 One card is drawn from a well-shuffled deck of 52 cards. Calculate the probability that the card will be an ace, Total number of cards = 52 Total number of aces = 4 P (getting an ace) = (𝑁𝑢𝑚𝑏𝑒𝑟 𝑜𝑓 𝑎𝑐𝑒𝑠)/(𝑇𝑜𝑡𝑎𝑙 𝑛𝑢𝑚𝑏𝑒𝑟 𝑜𝑓 𝑐𝑎𝑟𝑑𝑠) = 4/52 = 𝟏/𝟏𝟑 Example 4 One card is drawn from a well-shuffled deck of 52 cards. Calculate the probability that the card will (ii) not be an ace. P(we do not get an ace) = 1 – P(get an ace) = 1 – 𝟏/𝟏𝟑 = (13 −1)/13 = 𝟏𝟐/𝟏𝟑

Davneet Singh's photo - Co-founder, Teachoo

Made by

Davneet Singh

Davneet Singh is an IIT Kanpur graduate and has been teaching for 16+ years. At Teachoo, he breaks down Maths, Science and Computer Science into simple steps so students understand concepts deeply and score with confidence.

Many students prefer Teachoo Black for a smooth, ad-free learning experience.